Suppose the average income of a consumer decrease from R18 000 to R12 000. As a result, the quantity of product A demanded by the consumer increases from 200 units to 280 units. Use the ARC formula to calculate and classify the income the elasticity of demand for product A.



1
Expert's answer
2021-09-20T12:01:41-0400

Solution:

Arc Income elasticity of demand = "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; Income}"

Q1 = 200

Q2 = 280

Y1 = 18,000

Y2 = 12,000

Arc Ed = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 }\div \frac{I_{2} -I_{1}}{(I_{2}+I_{1})/2 }"


Arc Ed = "\frac{80}{240}\div \frac{-6,000}{15,000} = \frac{0.33}{-0.4} = -0.83"


Arc Income elasticity of demand = -0.83


Arc Income elasticity of demand for product A is negative which means that product A is an inferior good since quantity demanded increases with a decrease in income.